# Bridging Cache-Friendliness and Concurrency: A Locality-Optimized In-Memory B-Skiplist

### v2

@@ Appendix A Pseudocode
1:SkipList $SL$ and a pair key $K$ value $V$
[…]
@@
8:  while current.next_header <=<=< = key do
@@
29:      if rank +1<=+1<=+ 1 < = current.num_elements then
@@
34:       rank $\leftarrow$ rank -- current.num_elements
@@
37:     current.insert(K,V) at rank$+$1
38:     if level $>$then
39:      current.insert_child(rank$+$1)
@@
46:     current.split(new_node, rank$+$1)
47:     if level $>$then
48:      current.insert_child(rank$+$1)
@@ Appendix B Deadlock Freedom
Deadlock-freedom. We will show that the proposed concurrent B-skiplist is deadlock-free. The most popular way to avoid deadlock is to ensure that locks are acquired in a total order, thereby avoiding circular wait. In the proposed top-down concurrency control protocol, there is a total ordering on the locks from left-to-right within a level, and then from top-to-bottom in the levels. That is, the lowest ordered node is the left $\infty$- ∞ sentinel at the highest level, and the highest ordered node is the right $\infty$ sentinel at the leaf level. The left $\infty$- ∞ sentinel at some level $\ell$ is ordered directly after the right $\infty$ sentinel at level $\ell+1$. The insert and find algorithms acquire locks according to this total ordering in a left-to-right and top-to-bottom fashion, thereby avoiding deadlock.
@@ Appendix C Zipfian Key Figures
Load A B C E 111555101010Workload Normalized throughput NHSFolly SLJava SLB-skiplist
@@
Load A B C E 00.50.50.51111.51.51.5111Workload Normalized performance B-skiplistB-treeMassTree

### v3

@@ Appendix A Pseudocode
1:SkipList SLSL and a pair key $K$ value $V$
2:$K$,$V$ inserted into SLSL
@@
8:  while current.next_header <=<= key do
@@
29:      if rank +1<=+1<= current.num_elements then
@@
34:       rank $\leftarrow$ rank - current.num_elements
@@
37:     current.insert(K,V) at rank++1
38:     if level >>then
39:      current.insert_child(rank++1)
@@
46:     current.split(new_node, rank++1)
47:     if level >>then
48:      current.insert_child(rank++1)
@@ Appendix B Deadlock Freedom
Deadlock-freedom. We will show that the proposed concurrent B-skiplist is deadlock-free. The most popular way to avoid deadlock is to ensure that locks are acquired in a total order, thereby avoiding circular wait. In the proposed top-down concurrency control protocol, there is a total ordering on the locks from left-to-right within a level, and then from top-to-bottom in the levels. That is, the lowest ordered node is the left -$\infty$ sentinel at the highest level, and the highest ordered node is the right $\infty$ sentinel at the leaf level. The left -$\infty$ sentinel at some level $\ell$ is ordered directly after the right $\infty$ sentinel at level $\ell+1$. The insert and find algorithms acquire locks according to this total ordering in a left-to-right and top-to-bottom fashion, thereby avoiding deadlock.
@@ Appendix C Zipfian Key Figures
Load A B C E 11551010Workload Normalized throughput NHSFolly SLJava SLB-skiplist
@@
Load A B C E 00.50.5111.51.511Workload Normalized performance B-skiplistB-treeMassTree
@@ Appendix D Data Tables
Appendix E Artifact Instructions
The B-skiplist implementation is available at https://github.com/Ratbuyer/bskip_artifact, including all the tested systems in Section 5. The README files in the repository contain compiling and running instructions.